OpenStudy (kabase1234):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ... an = 4 • -3n + 1 an = -3 • 4n - 1 an = -3 • 4n an = 4 • -3n

OpenStudy (kabase1234):

@yanasidlinskiy

OpenStudy (kabase1234):

equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively. an = 7 • (-3)n + 1 an = 7 • 3n - 1 an = 7 • (-3)n - 1 an = 7 • 3n

OpenStudy (kabase1234):

@yanasidlinskiy

OpenStudy (sshayer):

\[a _{n}=ar ^{n-1}\] a=first term r=common ratio

OpenStudy (sshayer):

\[r=\frac{ a _{2} }{ a _{1} }=\frac{ -12 }{ -3 }=4\] a=-3 \[a _{n}=-3*4^{n-1}\]

OpenStudy (sshayer):

for second \[a _{n}=ar ^{n-1}\] \[a _{2}=ar ^{2-1}=ar=-21 ...(1)\] \[a _{5}=ar ^{5-1}=ar^4=567 ...(2)\] divide (2) by (1) \[\frac{ ar^4 }{ ar }=\frac{ 567 }{ -21 }\] \[r^3=-27,r=-3\] from (1) a=7 \[a _{n}=?\]